Michael is 6 years older than Vanessa. Seven years ago, Michael was 3 times as old as Vanessa. How old is Vanessa now?
Answer by htmentor(1343)

Let M = Michael's age now
Let V = Vanessa's age now
Right now Michael is 6 yrs older than Vanessa:
M = V + 6
7 years ago, Michael was 3 times as old as Vanessa:
M - 7 = 3(V-7)
(V+6) - 7 = 3V - 21
Solve for V:
V - 1 = 3V - 21
2V = 20
V = 10
So Vanessa is 10 yrs old and Michael is 16 yrs old.